Jack's construction Co. has 80,000 bonds outstanding that are selling at par value. Bonds with similar characteristics are yielding 8.5%. The company alson has 4 million shares of common stock outstanding. the stock has a beta of 1.1 and sells for $40 a share. The U.S. Treasury bill is yielding 4% and the market risk premium is 8%. Jack's tax rate is 35%. What is Jack's weighted average cost of capital?
Answer #1

\(\mathrm{R}_{\mathrm{e}}=.04+(1.1 \times .08)=.128\)

Debt: \(\quad 80,000 \times \$ 1,000=\$ 80 \mathrm{~m}\)

Common: \(\quad 4 \mathrm{~m} \times \$ 40=\$ 160 \mathrm{~m}\)

Total \(=\$ 80 \mathrm{~m}+\$ 160 \mathrm{~m}=\$ 240 \mathrm{~m}\)

\(\operatorname{WACC}=\left(\frac{\$ 160 \mathrm{~m}}{\$ 240 \mathrm{~m}} \times .128\right)+\left(\frac{\$ 80 \mathrm{~m}}{\$ 240 \mathrm{~m}} \times .085 \times(1-.35)\right)=.085333+.018417=.10375=\)

\(10.38\) percent